What is the impulse for a force of 40 newtons being applied to an object for 5 seconds?


1
Expert's answer
2021-06-07T09:29:20-0400

"F=\frac{\Delta p}{\Delta t}\to \Delta p= F\cdot \Delta t=40\cdot5=200\ (N\cdot s)" . Answer